OpenStudy (anonymous):

Find the derivative of f(x) = negative 9 divided by x at x = 6.

OpenStudy (p0sitr0n):

f=-9/x f' = 9/x^2 f'(6)=9/36 use the quotient rule, but usually [1/x]'=-1/x^2

OpenStudy (anonymous):

@P0sitr0n 2 divided by 3 1 divided by 4 3 divided by 2 4 these are my answer choices

OpenStudy (p0sitr0n):

im technically not supposed to give the answer right away, but i can tell that maybe theyre asking for f(x)? because 9/6 would be 3/2